One of the most important buffering systems in the biological realm is the carbonic acid (H2CO3)/carbonate ion (HCO3- ) system that maintains the pH of blood plasma to a relatively constant value. In blood at 37°C, the Ka of carbonic acid is 8x10-8. Suppose that the concentrations of the buffer components in the blood plasma are: [HCO3-] = 0.24 M and [H2CO3] = .12 M. Calculate the pH of blood.
Answer #1

using Buffer formula Prea pica + Log (ties) , pka = -log(+84168) = 7.1 = 7.1 + log ( 874 ) = 7.1+0.3 = 7.4 PH = 7.4